Let W = {(X1, X2, X3) €R³: X2 + X3 = 0}. How do I show that W is a subspace of R³? What are two subspaces W1 and W2 of R³ such that R³=W⊕W1and R³=W⊕W2 but W1≠W2?

1
Expert's answer
2021-07-21T13:14:57-0400

Let "p,q\in \mathbb{R}" and "(a_1,a_2,a_3),(b_1,b_2,b_3)\in W".

Consider

"p(a_1,a_2,a_3)+q(b_1,b_2,b_3)\\=(pa_1+qb_1,pa_2+qb_2,pa_3+qb_3)"

Now

"pa_2+qb_2+pa_3+qb_3\\=p(a_2+a_3)+q(b_2+b_3)=0"

So "p(a_1,a_2,a_3)+q(b_1,b_2,b_3)\in W", and hence "W" is a subspace of "\mathbb{R}^3".


Next, let "W_1=\{(x_2,x_2,0)\}" and "W_2=\{(x_2,2x_2,0)\}".

Suppose

"\textbf{s}\in W\cap W_1"

Then every element of "\textbf{s}" is of the form "(x_1,x_2,x_3)" where "x_3=0", "x_2=-x_3=0" and "x_1=x_2=0". Hence "\textbf{s}=\{(0,0,0)\}".

Also, if "(a,b,c)\in \mathbb{R}^3" then it can be written as

"(a,b,c)=(a-b-c,-c,c)+(b+c,b+c,0)"

So "\mathbb{R}^3=W\bigoplus W_1".

Similarly, suppose "\textbf{t}\in W \cap W_2". Then every element of "\textbf{t}" is of the form "(y_1,y_2,y_3)" where "y_3=0", "y_2=-y_3=0" and "y_1=y_2=0".

Also, if "(a,b,c)\in \mathbb{R}^3" then it can be written as

"(a,b,c)=\left(\frac{2a-b-c}{2},-c,c\right)+\left(\frac{b+c}{2},b+c,0\right)"

So "\mathbb{R}^3=W\bigoplus W_2".

Clearly, "W_1\neq W_2".